# Modeling Cell Dynamics and Interactions with Unbalanced Mean Field Schrödinger Bridge

### v2

@@ B.2 Mouse Blood Hematopoiesis
We obtained the dataset from (Weinreb et al., 2020). We used the cells profiled from in vitro cultured mouse hematopoietic progenitors, grown in conditions supporting multi-lineage differentiation. While not transplanted, the system retains extensive intercellular signaling, making it suitable for studying interaction-driven transition dynamics. The original gene expression space was projected to a 50-dimensional subspace using PCA. We summarized the results of hold-one-out evaluations in Table 7. The improvements observed in both Table 2 and Table 7 indicate that incorporating interaction terms indeed helps to recover the trajectories of cells. Moreover, the learned interacting forces exhibit significant correlation with the learned velocity, as quantified by a Moran’s I of 0.757. Based on our preceding analysis, this observed pattern and the improvement in performance strongly suggest the presence of genuine interaction forces within this realworld data.
@@
SF2M (Tong et al., 2024b) 12.863$\pm$0.004
Meta FM (Atanackovic et al., 2025) 14.079$\pm$0.000
MMFM (Rohbeck et al., 2025) 9.496$\pm$0.000
MIOFlow (Huguet et al., 2022) 7.779$\pm$0.000
uAM (Neklyudov et al., 2023) 9.157$\pm$0.000
TIGON (Sha et al., 2024) 8.402$\pm$0.000
DeepRUOT (Zhang et al., 2025a) 6.868$\pm$0.003
CytoBridge (Ours) 6.847$\pm$0.003
@@ B.3 Embryoid Body
We use the same dataset in (Moon et al., 2019; Tong et al., 2020), which consists of 16,819 cells collected at five time points. The cells are from human embryoid bodies (EBs), formed in 3D culture by spontaneous aggregation of stem cells over a 27-day differentiation time course. The experimental setup mimics early developmental conditions, where diverse cell types emerge and are expected to interact through signaling and spatial organization. We projected the original gene expression space to 50 dimensions using PCA. Hold-out evaluations were performed at four distinct time points, and our method consistently yielded the best results in these tests (Table 9). However, the observed performance gain, while positive, was less pronounced compared to the results on the mouse hematopoiesis data. Correspondingly, the learned interacting forces for this dataset exhibit a correlation pattern in the cell-state landscape with a Moran’s I of 0.450. This moderate positive Moran’s I suggests the presence of cell-state transition-related interaction effects within this data, though potentially weaker or less organized than those inferred for the lineage-constrained systems such as mouse hematopoiesis data.
@@
SF2M (Tong et al., 2024b) 9.146$\pm$0.003 0.748$\pm$0.000 11.398$\pm$0.018 0.377$\pm$0.000 14.288$\pm$0.054 0.539$\pm$0.000 19.804$\pm$0.112 0.399$\pm$0.000
Meta FM (Atanackovic et al., 2025) 9.531$\pm$0.000 0.748$\pm$0.000 12.443$\pm$0.000 0.377$\pm$0.000 13.928$\pm$0.000 0.539$\pm$0.000 15.903$\pm$0.000 0.399$\pm$0.000
MMFM (Rohbeck et al., 2025) 9.640$\pm$0.000 0.748$\pm$0.000 11.344$\pm$0.000 0.377$\pm$0.000 11.856$\pm$0.000 0.539$\pm$0.000 11.608$\pm$0.000 0.399$\pm$0.000
MIOFlow (Huguet et al., 2022) 8.447$\pm$0.000 0.748$\pm$0.000 9.229$\pm$0.000 0.377$\pm$0.000 9.436$\pm$0.000 0.539$\pm$0.000 10.123$\pm$0.000 0.399$\pm$0.000
uAM (Neklyudov et al., 2023) 12.315$\pm$0.000 1.486$\pm$0.000 14.996$\pm$0.000 1.323$\pm$0.000 15.685$\pm$0.000 1.526$\pm$0.000 18.407$\pm$0.000 1.396$\pm$0.000
UDSB (Bunne et al., 2023) 11.983$\pm$0.022 0.429$\pm$0.008 14.009$\pm$0.011 0.005$\pm$0.004 14.656$\pm$0.018 0.166$\pm$0.006 15.884$\pm$0.012 0.029$\pm$0.007
TIGON (Sha et al., 2024) 8.433$\pm$0.000 0.118$\pm$0.000 9.275$\pm$0.000 0.030$\pm$0.000 9.802$\pm$0.000 0.276$\pm$0.000 10.148$\pm$0.000 0.141$\pm$0.000
DeepRUOT (Zhang et al., 2025a) 8.159$\pm$0.002 0.050$\pm$0.001 9.034$\pm$0.003 0.161$\pm$0.002 9.369$\pm$0.003 0.005$\pm$0.005 9.773$\pm$0.007 0.262$\pm$0.007
CytoBridge (Ours) 8.159$\pm$0.002 0.002$\pm$0.001 9.027$\pm$0.003 0.057$\pm$0.002 9.351$\pm$0.005 0.175$\pm$0.007 9.750$\pm$0.006 0.022$\pm$0.006
@@
SF2M (Tong et al., 2024b) 12.257$\pm$0.002 13.603$\pm$0.013 13.567$\pm$0.002
Meta FM (Atanackovic et al., 2025) 16.523$\pm$0.000 16.417$\pm$0.000 15.019$\pm$0.000
MMFM (Rohbeck et al., 2025) 10.872$\pm$0.000 11.497$\pm$0.000 11.930$\pm$0.000
MIOFlow (Huguet et al., 2022) 10.684$\pm$0.000 11.755$\pm$0.000 10.440$\pm$0.000
uAM (Neklyudov et al., 2023) 12.857$\pm$0.000 15.743$\pm$0.000 17.433$\pm$0.000
TIGON (Sha et al., 2024) 11.199$\pm$0.000 11.207$\pm$0.000 10.833$\pm$0.000
DeepRUOT (Zhang et al., 2025a) 9.628$\pm$0.001 10.382$\pm$0.004 10.215$\pm$0.007
CytoBridge (Ours) 9.626$\pm$0.001 10.333$\pm$0.004 10.201$\pm$0.007
Figure 6: Application in embryoid body data ($\sigma=0.1$). (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=4$. (d) The interactions inferred by CytoBridge.
@@ B.4 Pancreatic $\beta$-cell Differentiation Data
We evaluated our method on the dataset from (Veres et al., 2019), which contains 51,274 cells collected across eight time points from human pluripotent stem cells differentiating toward pancreatic $\beta$-like cells in 3D suspension culture. The original gene expression space was projected to 30 dimensions using PCA. As shown in Fig. 7, CytoBridge is able to infer the velocity and growth rate, while identifying attractors. When comparing our approach, which explicitly models cellular interactions, to DeepRUOT, an algorithm that does not, we observed improved performance in 5 out of the 7 tested time points. Furthermore, the learned interacting forces for this dataset exhibited a moderate correlation distribution pattern with inferred state-transition velocity, quantified by a Moran’s I of 0.590. Overall, the performance gains across most tested time points, coupled with the observed moderate patterns in the inferred forces, may suggest the presence of interaction forces within this dataset. The correlation between forces and velocities indicates that cells may tend to prevent others from differentiation at early stages while promoting differentiation at later time points.
@@
DeepRUOT 8.0447$\pm$0.0005 8.0773$\pm$0.0021 7.6301$\pm$0.0032 8.0064$\pm$0.0042 7.9018$\pm$0.0117 8.3977$\pm$0.0102 7.8346$\pm$0.0109
CytoBridge (Ours) 8.0448$\pm$0.0005 8.0771$\pm$0.0021 7.6299$\pm$0.0032 8.0066$\pm$0.0043 7.9018$\pm$0.0117 8.3974$\pm$0.0102 7.8343$\pm$0.0109
Figure 7: Application in pancreatic $\beta$-cell differentiation data ($\sigma=0.1$). (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=7$. (d) The interactions inferred by CytoBridge.
@@ B.5 EMT Data
We use the dataset from (Sha et al., 2024; Cook & Vanderhyden, 2020), derived from A549 cancer cells undergoing TGFB1-induced epithelial-mesenchymal transition (EMT). This dataset comprises four distinct time points. The cells were cultured in standard 2D monolayers and treated with TGFB1 over several days. Although EMT is a coordinated process in vivo involving paracrine and contact-dependent signaling, this in vitro system mimics EMT as a largely cell-autonomous response to an external stimulus. Despite the fact that CytoBridge still is able to infer the velocities, the growth of cells, and different cell fates (Fig. 8). When comparing our method, which explicitly models interaction terms, to approaches like DeepRUOT that do not, we observed no improvement in performance on this dataset. Furthermore, the distribution pattern of the learned interacting forces for this dataset was very disorganized, yielding a Moran’s I of 0.040. This indicates that the inferred forces are largely random compared to the transition velocity direction, suggesting a lack of significant or organized intercellular interactions for state-transition. The absence of both performance gain and significant correlation in inferred forces suggests that transitions in this setting are primarily driven by direct transcriptional responses rather than intercellular signaling, which is consistent with the biological experiment setup.
@@
Figure 8: Application in EMT data ($\sigma=0.05$). (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=4$. (d) The interactions inferred by CytoBridge.
Figure 9: The correlation of drift and interacting forces inferred from (a) mouse blood hematopoiesis (b) embryoid body (c) pancreatic $\beta$-cell differentiation data (d) EMT data.
B.6 Ablation Studies
@@
CytoBridge w/o interaction (DeepRUOT) 0.044$\pm$0.002 0.014$\pm$0.007 0.045$\pm$0.002 0.026$\pm$0.018 0.053$\pm$0.002 0.059$\pm$0.032 0.057$\pm$0.003 0.075$\pm$0.044
CytoBridge w/o growth 0.068$\pm$0.001 0.080$\pm$0.000 0.175$\pm$0.003 0.250$\pm$0.000 0.343$\pm$0.010 0.515$\pm$0.000 0.480$\pm$0.018 0.930$\pm$0.000
CytoBridge w/o $\mathcal{L}_{\text{mass}}$ 0.016$\pm$0.001 0.042$\pm$0.014 0.037$\pm$0.004 0.165$\pm$0.030 0.046$\pm$0.007 0.352$\pm$0.041 0.047$\pm$0.005 0.689$\pm$0.046
CytoBridge w/o $\mathcal{L}_{\text{FP}}$ 0.014$\pm$0.000 0.013$\pm$0.012 0.016$\pm$0.001 0.030$\pm$0.019 0.022$\pm$0.003 0.048$\pm$0.038 0.039$\pm$0.004 0.063$\pm$0.058
CytoBridge w/o pretraining 0.426$\pm$0.003 0.016$\pm$0.013 0.915$\pm$0.003 0.035$\pm$0.016 1.195$\pm$0.003 0.044$\pm$0.040 1.451$\pm$0.004 0.168$\pm$0.081
CytoBridge w/o score matching 0.025$\pm$0.001 0.011$\pm$0.011 0.025$\pm$0.002 0.024$\pm$0.028 0.034$\pm$0.003 0.049$\pm$0.039 0.039$\pm$0.003 0.084$\pm$0.051
CytoBridge w/o training 0.036$\pm$0.002 0.012$\pm$0.006 0.029$\pm$0.003 0.021$\pm$0.025 0.067$\pm$0.003 0.050$\pm$0.025 0.108$\pm$0.002 0.174$\pm$0.076
CytoBridge (Ours) 0.015$\pm$0.001 0.013$\pm$0.009 0.014$\pm$0.001 0.021$\pm$0.024 0.018$\pm$0.002 0.043$\pm$0.041 0.038$\pm$0.003 0.058$\pm$0.061
@@ C.1 Evaluation Metrics
To evaluate the performance of SF2M (Tong et al., 2024b), we keep the diffusion coefficient $\sigma$ the same as ours while maintaining other parameters as defaults for fair comparison. To evaluate the performance of Meta FM (Atanackovic et al., 2025), Multi-Marginal FM (MMFM) (Rohbeck et al., 2025), and MIOFlow (Huguet et al., 2022), we use their default parameters. The weights of the inferred particles are set to uniform distribution to evaluate as the growth term is not considered. To evaluate the performance of UDSB (Pariset et al., 2023), as its default setting only involves three-time points, we use samples at $t=0,2,4$ from synthetic gene data and embryoid body data as inputs.
@@ C.2 Hyperparameters Selection and Loss Weighting
The experiments were performed on a shared high-performance computing cluster with NVIDIA A100 GPU and 128 CPU cores. As we aim to make our algorithm universally applicable to different types of biological data, most of the hyperparameters are kept identical across different datasets, while those varying are mainly related to the scale of datasets. Specifically, for the modeling of interaction potential, we set the number of RBF kernels to 8 across different datasets. For real-world scRNA-seq data, the threshold $d_{\text{cutoff}}$ is set no lower than the largest distance between cells in specific datasets so that all pairs of cells are involved in interacting with each other. We also conducted experiments on the effect of different values of $d_{\text{cutoff}}$. As shown in Table 13, we found that the performance was quite robust with respect to different choices of $d_{\text{cutoff}}$ unless the cutoff is set too small, which may potentially lead to inadequate modeling of cellular interactions. To simulate ODEs with random batch methods, we set the number of particles $p$ within one group to 16. For our training procedure, the parameters differ only in the number of training epochs. As real-world scRNA-seq data mainly involves large numbers of cells, it typically will require more iterations for our model to converge.
@@
0.1 0.080$\pm$0.001 0.010$\pm$0.008 0.119$\pm$0.003 0.032$\pm$0.016 0.141$\pm$0.003 0.050$\pm$0.042 0.167$\pm$0.003 0.077$\pm$0.043
0.3 0.015$\pm$0.001 0.013$\pm$0.011 0.016$\pm$0.001 0.024$\pm$0.023 0.026$\pm$0.003 0.038$\pm$0.044 0.042$\pm$0.004 0.052$\pm$0.063
0.5 0.015$\pm$0.001 0.013$\pm$0.009 0.014$\pm$0.001 0.021$\pm$0.024 0.018$\pm$0.002 0.043$\pm$0.041 0.038$\pm$0.003 0.058$\pm$0.061
0.7 0.013$\pm$0.001 0.013$\pm$0.007 0.013$\pm$0.001 0.024$\pm$0.023 0.024$\pm$0.003 0.047$\pm$0.031 0.037$\pm$0.003 0.069$\pm$0.046
1.0 0.014$\pm$0.001 0.012$\pm$0.009 0.014$\pm$0.001 0.025$\pm$0.023 0.026$\pm$0.004 0.039$\pm$0.041 0.043$\pm$0.004 0.057$\pm$0.056
2.0 0.018$\pm$0.001 0.014$\pm$0.011 0.025$\pm$0.002 0.023$\pm$0.021 0.042$\pm$0.003 0.046$\pm$0.031 0.049$\pm$0.005 0.072$\pm$0.037
@@
For other hyper-parameters, we will then provide a guideline on their choices. In the pre-training phase, we first need to provide a suitable initialization for the velocity network and the growth network, which involves selecting the parameters $\lambda_{m}$ and $\lambda_{d}$ in $\mathcal{L}_{\text{Recons}}=\lambda_{m}\mathcal{L}_{\text{Mass}}+\lambda_{d}\mathcal{L}_{\text{OT}}$. We set $\lambda_{d}$ to 1 and $\lambda_{m}$ to 0.01 in order to encourage the transition to match the observed distribution while maintaining the unbalanced effect. Here we only adopt the local mass matching loss without restricting the exact number of cells in order to learn general growth patterns. Subsequently, we set $\lambda_{m}$ to zero to initialize the interaction network. By doing so, the interaction network can be stably trained in order to refine the variance of distributions. The parameter selections of pre-training procedure is summarized in Table 14, where the arrow notation ($\rightarrow$) represents the adjustment of hyperparameters during two stages of our pre-training phase.
@@
During the training phase, as the four networks have been reasonably initialized, these networks can be trained together stably. We set $\lambda_{m}$ and $\lambda_{d}$ to 1, while adding the global mass matching term to encourage our growth network to match the exact number of cells at different time points. $\lambda_{r}$ in $\mathcal{L}=\mathcal{L}_{\text{Energy}}+\lambda_{r}\mathcal{L}_{\text{Recons}}+\lambda_{f}\mathcal{L}_{\text{FP}}$ is set to better match the distributions. $\lambda_{f}$ and $\lambda_{w}$ in $\mathcal{L}_{\text{FP}}$ are set to align the score network to match the density. The specific choices of these parameters are summarized in Table 14. We utilize the set of parameters consistently across different datasets.
@@
Table 14: Parameter Settings for Different Datasets Across Two Training Stages (Synthetic gene, mouse hematopoiesis, embryoid body, pancreatic $\beta$-cell differentiation, EMT).
@@
Parameter Synthetic gene mouse hematopoiesis embryoid body pancreatic $\beta$-cell differentiation EMT
@@
$(\lambda_{m},\lambda_{d},\text{Epochs})$ (0.01, 1, 200) $\rightarrow$ (0.0, 1, 100) (0.01, 1, 500) $\rightarrow$ (0.0, 1, 100) (0.01, 1, 500) $\rightarrow$ (0.0, 1, 100) (0.01, 1, 500) $\rightarrow$ (0.0, 1, 100) (0.01, 1, 100) $\rightarrow$ (0.0, 1, 100)

### v3

@@ B.2 Mouse Blood Hematopoiesis
We obtained the dataset from (Weinreb et al., 2020). We used the cells profiled from in vitro cultured mouse hematopoietic progenitors, grown in conditions supporting multi-lineage differentiation. While not transplanted, the system retains extensive intercellular signaling, making it suitable for studying interaction-driven transition dynamics. The original gene expression space was projected to a 50-dimensional subspace using PCA. We summarized the results of hold-one-out evaluations in Table˜7. The improvements observed in both Table˜2 and Table˜7 indicate that incorporating interaction terms indeed helps to recover the trajectories of cells. Moreover, the learned interacting forces exhibit significant correlation with the learned velocity, as quantified by a Moran’s I of 0.757. Based on our preceding analysis, this observed pattern and the improvement in performance strongly suggest the presence of genuine interaction forces within this realworld data. To enhance interpretability, we identified top 200 genes influenced most significantly by cell-cell interactions. Subsequent enrichment analysis of these genes revealed key pathways associated with the biological processes, as shown in Table˜8. Applying this to the mouse hematopoiesis dataset, we identified pathways related to positive regulation of leukocyte activation and cell-cell adhesion which are closely linked to hematopoiesis, interactions, and differentiation. We also incorporated Trajectory Inference with Cell–Cell Interactions (TICCI) (Fu et al., 2025) as a reference method. By applying TICCI to the mouse hematopoietic dataset, it identified ligand-recptor pair Lgals9-Cd45, which is closely related to the regulation of T cell activation, further provides biological interpretation for enriched pathways. Moreover, we calculated the gradients of growth network with respect to genes to identify key genes that contribute to growth dynamics, including Meis1, Nfkb1, and Rap1b. Meis1 regulates hematopoietic stem cell proliferation and self-renewal, Nfkb1 promotes cell cycle entry and survival as a signaling hub, and Rap1b drives cell division via pathways like MAPK, consistent with established biological knowledge. These findings demonstrate CytoBridge’s ability to uncover biologically relevant mechanisms.
@@
[…]
Table 8: Enriched pathways from interactions of mouse hematopoiesis data, showing the adjusted p-value and gene count for each term.
Pathway p.adjust Count
regulation of T cell activation $3.38\times 10^{-12}$ 23
positive regulation of cell activation $1.42\times 10^{-11}$ 23
positive regulation of cell-cell adhesion $5.81\times 10^{-11}$ 20
positive regulation of leukocyte activation $1.90\times 10^{-10}$ 21
leukocyte cell-cell adhesion $3.18\times 10^{-10}$ 21
lymphocyte differentiation $1.04\times 10^{-9}$ 21
@@ B.3 Embryoid Body
We use the same dataset in (Moon et al., 2019; Tong et al., 2020), which consists of 16,819 cells collected at five time points. The cells are from human embryoid bodies (EBs), formed in 3D culture by spontaneous aggregation of stem cells over a 27-day differentiation time course. The experimental setup mimics early developmental conditions, where diverse cell types emerge and are expected to interact through signaling and spatial organization. We projected the original gene expression space to 50 dimensions using PCA. Hold-out evaluations were performed at four distinct time points, and our method consistently yielded the best results in these tests (Table˜10). However, the observed performance gain, while positive, was less pronounced compared to the results on the mouse hematopoiesis data. Correspondingly, the learned interacting forces for this dataset exhibit a correlation pattern in the cell-state landscape with a Moran’s I of 0.450. This moderate positive Moran’s I suggests the presence of cell-state transition-related interaction effects within this data, though potentially weaker or less organized than those inferred for the lineage-constrained systems such as mouse hematopoiesis data.
@@
[…]
@@
[…]
Figure 6: Application in embryoid body data ($\sigma=0.1$), visualized in PHATE space. (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=4$. (d) The correlation of velocity and interacting forces.
@@ B.4 Pancreatic $\beta$-cell Differentiation Data
We evaluated our method on the dataset from (Veres et al., 2019), which contains 51,274 cells collected across eight time points from human pluripotent stem cells differentiating toward pancreatic $\beta$-like cells in 3D suspension culture. The original gene expression space was projected to 30 dimensions using PCA. As shown in Fig.˜7, CytoBridge is able to infer the velocity and growth rate, while identifying attractors. When comparing our approach, which explicitly models cellular interactions, to DeepRUOT, an algorithm that does not, we observed improved performance in 5 out of the 7 tested time points. Furthermore, the learned interacting forces for this dataset exhibited a moderate correlation distribution pattern with inferred state-transition velocity, quantified by a Moran’s I of 0.590. Overall, the performance gains across most tested time points, coupled with the observed moderate patterns in the inferred forces, may suggest the presence of interaction forces within this dataset. The correlation between forces and velocities indicates that cells may tend to prevent others from differentiation at early stages while promoting differentiation at later time points.
@@
[…]
Figure 7: Application in pancreatic $\beta$-cell differentiation data ($\sigma=0.1$), visualized in UMAP space. (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=7$. (d) The correlation of velocity and interacting forces.
@@ B.5 EMT Data
We use the dataset from (Sha et al., 2024; Cook and Vanderhyden, 2020), derived from A549 cancer cells undergoing TGFB1-induced epithelial-mesenchymal transition (EMT). This dataset comprises four distinct time points. The cells were cultured in standard 2D monolayers and treated with TGFB1 over several days. Although EMT is a coordinated process in vivo involving paracrine and contact-dependent signaling, this in vitro system mimics EMT as a largely cell-autonomous response to an external stimulus. Despite the fact that CytoBridge still is able to infer the velocities, the growth of cells, and different cell fates (Fig.˜8). When comparing our method, which explicitly models interaction terms, to approaches like DeepRUOT that do not, we observed no improvement in performance on this dataset. Furthermore, the distribution pattern of the learned interacting forces for this dataset was very disorganized, yielding a Moran’s I of 0.040. This indicates that the inferred forces are largely random compared to the transition velocity direction, suggesting a lack of significant or organized intercellular interactions for state-transition. The absence of both performance gain and significant correlation in inferred forces suggests that transitions in this setting are primarily driven by direct transcriptional responses rather than intercellular signaling, which is consistent with the biological experiment setup.
@@
Figure 8: Application in EMT data ($\sigma=0.05$), visualized in PCA space. (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=4$. (d) The correlation of velocity and interacting forces.
B.6 Zebrafish Spatiotemporal Data
We adopt the Zebrafish Embryogenesis Spatiotemporal Transcriptomic Atlas (ZESTA) (Liu et al., 2022a), created using the Stereo-seq spatial transcriptomics technology (Chen et al., 2022a). The dataset provides a high-resolution map of gene expression in zebrafish embryos across six critical time points within the first 24 hours of development. Among the six time points, we selected 5.25 hpf and 10 hpf as input. To align spatial coordinates between these two time points, we adopt the rigid body transformation invariant optimal transport. Then, we projected the original gene expression space to 50 dimensions using PCA. To model cellular interactions in both physical space and gene expression space, we transformed both the distances in physical and gene expression space into separate RBF features, and concatenated them as inputs to the interaction potential. Thus, the effects of cellular interactions are achieved by calculating the gradients of interaction potential with respect to distances in physical and gene expression space respectively. The reconstruction loss is calculated in both physical space and gene expression space. We compared the $\mathcal{W}_{1}$ distances in the physical space and gene expression space between CytoBridge and other methods. As shown in Table˜13, CytoBridge achieves better performance over other methods in both physical space and gene expression space. We also visualize the predicted cell states in physical space and gene expression space in Fig.˜9. To further interpret the biological effects of learned cellular interactions on the development process of zebrafish, we identified top 200 genes influenced most significantly by interactions. We identified pathways related to somite development which are critical to zebrafish embryonic development based on the enrichment analysis, as shown in Table˜14. These findings align with known biological processes, showing the framework’s potential in spatially resolved data.
Table 13: Wasserstein distance ($\mathcal{W}_{1}$) of predictions on zebrafish embryogenesis data. We report metrics in physical space (denoted as ’Space’) and gene expression space (denoted as ’Gene’). Bold indicates the best result.
Model Space Gene
SF2M (Tong et al., 2024b) 0.265 5.423
Meta FM (Atanackovic et al., 2025) 0.268 5.413
MMFM (Rohbeck et al., 2025) 0.247 5.208
Metric FM (Kapusniak et al., 2024) 0.273 5.366
UOT-FM (Eyring et al., 2024) 0.227 5.173
MIOflow (Huguet et al., 2022) 0.263 4.720
uAM (Neklyudov et al., 2023) 0.177 6.499
TIGON (Sha et al., 2024) 0.352 4.979
DeepRUOT (Zhang et al., 2025a) 0.261 4.745
CytoBridge (Ours) 0.035 4.712
Table 14: Enriched pathways from zebrafish embryogenesis data, showing the adjusted p-value and gene count for each term.
Pathway p.adjust Count
nucleolus $5.76\times 10^{-5}$ 12
somite development $1.82\times 10^{-4}$ 10
lipid transport $1.82\times 10^{-4}$ 12
gastrulation $4.15\times 10^{-4}$ 12
somitogenesis $4.15\times 10^{-4}$ 8
Figure 9: Application in zebrafish embryogenesis data ($\sigma=0.05$). (a) The spatial coordinates of cells sampled at $T=0$. (b) The predicted spatial coordinates of cells at $T=1$ by CytoBridge (c) The predicted gene expression by CytoBridge in the PCA space. (d) The overall velocity learned by CytoBridge in gene expression space.
B.7 Ablation Studies
@@
[…]
@@ C.1 Evaluation Metrics
To ensure a fair comparison with other methods, we used their default settings for datasets featured in their original papers. For all other datasets, we adjusted the models’ network sizes to ensure comparable parameter counts and tuned their training epochs and learning rates accordingly. To evaluate the performance of SF2M (Tong et al., 2024b), we keep the diffusion coefficient $\sigma$ the same as ours while maintaining other parameters as defaults for fair comparison. The weights of the inferred particles are set to uniform distribution to evaluate as the growth term is not considered. To evaluate the performance of UDSB (Pariset et al., 2023), as its default setting only involves three-time points, we use samples at $t=0,2,4$ from synthetic gene data and embryoid body data as inputs.
@@ C.2 Hyperparameters Selection and Loss Weighting
The experiments were performed on a shared high-performance computing cluster with NVIDIA A100 GPU and 128 CPU cores. As we aim to make our algorithm universally applicable to different types of biological data, most of the hyperparameters are kept identical across different datasets, while those varying are mainly related to the scale of datasets. Specifically, for the modeling of interaction potential, we set the number of RBF kernels to 8 across different datasets. For real-world scRNA-seq data, the threshold $d_{\text{cutoff}}$ is set no lower than the largest distance between cells in specific datasets so that all pairs of cells are involved in interacting with each other. For the zebrafish data, we set the threshold $d_{\text{cutoff}}$ in the physical space, where cells interact with neighboring cells in physical space. We conducted experiments on the effect of different values of $d_{\text{cutoff}}$ in Table˜16.
@@
[…]
As shown, we found that the performance was quite robust with respect to different choices of $d_{\text{cutoff}}$ unless the cutoff is set too small, which may potentially lead to inadequate modeling of cellular interactions. To simulate ODEs with random batch methods, we also examined the choices of number of particles $p$ within one group in Table˜17. Generally, increasing $p$ slightly improves performance on evaluation metrics by providing a more accurate approximation of the interaction term. However, gains diminish at higher $p$ values, while computational costs increase. Thus, we selected $p=16$ as the default, balancing robust performance with computational efficiency.
Table 17: Wasserstein distance ($\mathcal{W}_{1}$) across different time points for varying values of parameter pp.
$p$ $t=1$ $t=2$ $t=3$ $t=4$
2 0.025 0.033 0.048 0.094
4 0.017 0.023 0.025 0.043
8 0.016 0.019 0.022 0.035
16 0.015 0.014 0.018 0.038
32 0.015 0.016 0.020 0.030
64 0.014 0.014 0.024 0.030
Moreover, we also compared the results with and without the random batch method in Table˜18. Results show that the full model (without RBM) and the RBM model yield comparable performance, with the full model slightly better at certain time points. However, RBM significantly enhances computational efficiency, reducing the interaction term’s complexity from $O(N^{2})$ to $O(pN)$. The tables demonstrate substantial reductions in memory and inference time with RBM. Given its comparable accuracy and scalability for large-scale single-cell datasets, RBM is a justified and essential component of our framework.
Table 18: Performance comparison on simulation data with and without RBM.
$t=1$ $t=2$ $t=3$ $t=4$ Time (s) Memory (GB)
Method $\mathcal{W}_{1}$ $\mathcal{W}_{1}$ $\mathcal{W}_{1}$ $\mathcal{W}_{1}$
w/o RBM 0.015 0.013 0.023 0.029 0.568 22.1
$p=16$ 0.015 0.014 0.018 0.038 0.115 1.7
@@
For our training procedure, the parameters differ only in the number of training epochs. As real-world scRNA-seq data mainly involves large numbers of cells, it typically will require more iterations for our model to converge. Increasing the number of training epochs has few adverse effect, allowing users to use a default of (500, 100, 500) epochs for larger datasets. For other hyper-parameters, we will then provide a guideline on their choices. In the pre-training phase, we first need to provide a suitable initialization for the velocity network and the growth network, which involves selecting the parameters $\lambda_{m}$ and $\lambda_{d}$ in $\mathcal{L}_{\text{Recons}}=\lambda_{m}\mathcal{L}_{\text{Mass}}+\lambda_{d}\mathcal{L}_{\text{OT}}$. We set $\lambda_{d}$ to 1 and $\lambda_{m}$ to 0.01 in order to encourage the transition to match the observed distribution while maintaining the unbalanced effect. We empirically found that lowering the mass loss weighting during pre-training improves distribution matching performance. Here we only adopt the local mass matching loss without restricting the exact number of cells in order to learn general growth patterns. Subsequently, we set $\lambda_{m}$ to zero to initialize the interaction network. By doing so, the interaction network can be stably trained in order to refine the variance of distributions. The parameter selections of pre-training procedure is summarized in Table˜19, where the arrow notation ($\rightarrow$) represents the adjustment of hyperparameters during two stages of our pre-training phase.
@@
During the training phase, as the four networks have been reasonably initialized, these networks can be trained together stably. We set $\lambda_{m}$ and $\lambda_{d}$ to 1, while adding the global mass matching term to encourage our growth network to match the exact number of cells at different time points. $\lambda_{r}$ in $\mathcal{L}=\mathcal{L}_{\text{Energy}}+\lambda_{r}\mathcal{L}_{\text{Recons}}+\lambda_{f}\mathcal{L}_{\text{FP}}$ is set to better match the distributions. $\lambda_{f}$ and $\lambda_{w}$ in $\mathcal{L}_{\text{FP}}$ are set to align the score network to match the density. The specific choices of these parameters are summarized in Table˜19. We utilize the set of parameters consistently across different datasets.
@@
Table 19: Parameter Settings for Different Datasets Across Two Training Stages (Synthetic gene, mouse hematopoiesis, embryoid body, pancreatic $\beta$-cell differentiation, EMT, Zebrafish).
@@
Parameter Synthetic gene mouse hematopoiesis embryoid body pancreatic $\beta$-cell differentiation EMT Zebrafish
@@
$(\lambda_{m},\lambda_{d},\text{Epochs})$ (0.01, 1, 200) (0.01, 1, 500) (0.01, 1, 500) (0.01, 1, 500) (0.01, 1, 100) (0.01, 1, 500)
$\rightarrow$ (0.0, 1, 100) $\rightarrow$ (0.0, 1, 100) $\rightarrow$ (0.0, 1, 100) $\rightarrow$ (0.0, 1, 100) $\rightarrow$ (0.0, 1, 100) $\rightarrow$ (0.0, 1, 100)